# A Mathematical Theory for Learning Semantic Languages by Abstract Learners

### v2

@@ Abstract
Recent advances in Large Language Models (LLMs) have demonstrated the emergence of capabilities (learned skills) when the number of system parameters and the size of training data surpass certain thresholds. The exact mechanisms behind such phenomena are not fully understood and remain a topic of active research. Inspired by the skill-text bipartite graph model proposed by Arora and Goyal for modeling semantic languages, we develop a mathematical theory to explain the emergence of learned skills, taking the learning (or training) process into account. Our approach models the learning process for skills in the skill-text bipartite graph as an iterative decoding process in Low-Density Parity Check (LDPC) codes and Irregular Repetition Slotted ALOHA (IRSA). Using density evolution analysis, we demonstrate the emergence of learned skills when the ratio of the size of training texts to the number of skills exceeds a certain threshold. Our analysis also yields a scaling law for testing errors relative to the size of training texts. Upon completion of the training, we propose a method for semantic compression and discuss its application in semantic communication.
@@ I Introduction
In the recent era of natural language processing (NLP), the evolution of large language models (LLMs), such as GPT-4 [1] and Gemini [2], has greatly impacted people’s daily lives. There is a growing consensus that enhancing language model performance and sample efficiency across a broad spectrum of downstream NLP tasks is closely linked to scaling up these models. This scaling involves increasing both the size of the training data and the number of model parameters, as discussed in [3, 4]. The relationship between the scale of an LLM and its performance can often be quantitatively forecasted using scaling laws, as explored in-depth in [5, 6].
@@
    Learning as an Iterative Decoding Process: We model the learning/training process for the skills in the skill-text bipartite graph as an iterative decoding process, similar to the analysis in Low-Density Parity Check (LDPC) codes [14, 15, 16] and Irregular Repetition Slotted ALOHA (IRSA) [17, 18, 19, 20, 21, 22]. When a text is presented to a learner, a skill in that text might be learned with a certain probability (specific definitions for learners are given in Section III). By repeatedly presenting a large number of training texts to a learner, a fraction of skills can be learned.
@@
    The Emergence of Learned Skills: Following the density evolution analysis for the iterative decoding process (in the asymptotic regime of a large number of skills) [23, 24, 25], we show the emergence of learned skills when the ratio $R$ of the number of training texts to the number of skills exceeds a certain threshold. Once this threshold is exceeded, the testing error, defined as the probability of whether the learner can understand a randomly selected text, drops sharply. Our analysis also provides the scaling law of the testing error with respect to the ratio $R$.
@@
    The Association of Skills: Upon completing training, a learner acquires not just isolated skills but also understands how these skills interrelate, which can be represented by a skill association graph where edges connect learned skills appearing in the same text. Understanding this graph is crucial for making inferences, such as predicting the next skill from a given set. This process is akin to prompting a language model to forecast and generate text based on a skill set. This leads to a critical inquiry: Are the learned skills sufficiently interconnected to aid in prediction? To address this question, we use site (or node) percolation analysis [9] to derive conditions for the existence of a giant component in the skill association graph.
@@
  • (vi)
The paper is organized as follows: In Section II, we define a semantic language using a skill-text bipartite graph. In Section III, we introduce abstract learners, including 1-skill learners in Section III-A and Poisson learners in Section III-B. Density evolution analysis is employed to derive the scaling law for these abstract learners. In Section IV, we address the problem of learning the association between two learned skills and derive conditions for the existence of a giant component in the skill association graph. In Section V, we further address the problem setting with a hierarchy of skills by using a foundation model and a fine-tuning model. In Section VI, we demonstrate how the trained learners can be utilized for semantic compression and semantic communication. We present the extension to the setting with multiple classes of skills and texts in Section VII. The paper concludes in Section VIII, where we discuss possible extensions of our work.
@@ II-A Definition of a Semantic Language
(Semantic Language) A semantic language ${\cal L}=({\cal A},T,S,\phi)$ consists of (i) a set of tokens (symbols) ${\cal A}$, (ii) a set of texts $T$ composed of sequences of tokens, (iii) a set of skills $S$, and (iv) a function $\phi$ that maps a text $t$ in $T$ to a set of skills.
Two texts $t_{1}$ and $t_{2}$ are said to be semantically equivalent (or simply equivalent) if they both require the same set of skills, i.e., $\phi(t_{1})=\phi(t_{2})$. We say that a skill $s$ is present in a text $t$ if $s$ is included in the set of skills $\phi(t)$. For a semantic language ${\cal L}=({\cal A},T,S,\phi)$, we can generate a skill-text bipartite graph $G=(S,T,E)$, where an edge is added between a skill $s$ and a text $t$ if $s$ is present in $\phi(t)$ (see Figure 1 for an illustration). Note that the skill-text bipartite graph $G=(S,T,E)$ characterizes the semantic language.
@@
In this paper, we address the question of learning a semantic language. A common approach is to sample a subset of texts from the set of texts $T$ and present this subset to a learner to learn the mapping $\phi$.
Definition 2
A skill $s$ in a semantic language is learned if the learner can determine whether the skill $s$ is present in any given text $t$ in $T$. A text $t$ is understood by a learner if all the skills contained within $t$ are learned by the learner.
II-B Learning a Semantic Language by Sampling
As mentioned in the previous section, we sample a subset of $D$ texts from the set of texts $T$ and present these $D$ texts as the training texts. Let $R=D/|S|$ be the ratio of the number of training texts to the number of skills.
We make the following two assumptions about a randomly selected text:
(Poisson distribution) For a randomly selected text $t\in T$, the number of skills in $t$ follows a Poisson random variable with mean $c$, i.e., ${\bf\sf P}(|\phi(t)|=k)=e^{-c}\frac{c^{k}}{k!}$ for $k=0,1,2,\ldots$.
(Uniform distribution) The probability that an edge of the (randomly selected) text node $t$ is connected to a skill node $s$ is $1/|S|$.
From (A1)-(A2), the skill-text bipartite graph of a sampled semantic language is a random bipartite graph with $|S|$ skill nodes on one side and $D$ text nodes on the other side. An edge exists between a skill node $s$ and a text node $t$ if $s\in\phi(t)$. The degree of a text node follows a Poisson distribution with mean $c$. According to (A1) and (A2), the number of edges between a text node and a particular skill node follows a Poisson distribution with mean $c/|S|$. Since there are $D$ text nodes, the number of edges of a skill node is also Poisson distributed, with mean $cD/|S|$ (i.e., $cR$). This leads to the following proposition.
@@ II-B Learning a Semantic Language by Sampling
Consider the sampled skill-text bipartite graph from a semantic language ${\cal L}=({\cal A},T,S,\phi)$. Then: (i) The degree of a skill node follows a Poisson distribution with mean $cD/|S|=cR$. (ii) The probability that a randomly selected edge is connected to a particular text node is $1/D$, indicating a uniform distribution. Therefore, such a bipartite graph is a random bipartite graph, where the degree distribution of the text nodes (respectively, skill nodes) is Poisson with mean $c$ (respectively, $cR$).
@@ III Abstract Learners
In this section, we develop a mathematical theory to understand the learning process of a semantic language by an abstract learner. To a learner, all skills are novel before training. After presenting a text $t$ to the learner, a skill $s$ in the text $t$ might be learned. After repeatedly presenting the training set of $D$ texts, we aim to determine the fraction of skills that are learned by the learner. We consider two types of learners based on their learning capabilities, namely, a 1-skill learner and a more general Poisson learner.
@@
(1-Skill Learner) An abstract learner is called a 1-skill learner if it can learn a novel skill $s$ by being presented with a text $t$ where the skill $s$ is the only novel skill in the text. Once a skill $s$ is learned, the learner is able to identify the skill $s$ appearing in any other texts.
@@
Now we describe how the 1-skill learner is trained. For each iteration, we present the $D$ sampled texts to the 1-skill learner in parallel. In the first iteration, a text containing only one skill is used to learn the skill in that text. Texts containing more than one skill do not contribute to learning in the first iteration. Once a skill $s$ is learned by the 1-skill learner, the number of novel skills in other texts containing $s$ is reduced by 1. In other words, the edges connected to the skill node $s$ can be removed from the skill-text bipartite graph. Therefore, the number of novel skills associated with each text in the next iteration will not be greater than that in the current iteration. In the second iteration, texts with only one novel skill remaining are used to learn the skills in those texts. As in the first iteration, skills learned in the second iteration can be used to remove the corresponding edges in the skill-text bipartite graph. This iteration process is repeated until no more novel skills can be learned. We refer to this training process as Successive Cancellation of Novel Skills (SCNS).
Figure 2: The Successive Cancellation of Novel Skills (SCNS) example.
@@
In Figure 2, we present a simple example of the SCNS. Figure 2(a) depicts the original skill-text bipartite graph. Figure 2(b) focuses on the skill node $a$, which is connected to a text node with degree $1$; hence, we use a red edge to indicate that the skill node $a$ is learned by the text node. Consequently, the skill node $a$ can be learned, allowing the edges connected to it to be removed from the skill-text bipartite graph. Figure 2(c) shows the graph after the edges are removed, revealing that the skill node $b$ is connected to a text node with degree $1$, indicating it can also be learned. Figure 2(d) represents the last iteration, where no additional skills can be learned, concluding the process with the skills $a$ and $b$ being learned, while skills $c$ and $d$ are not.
@@
This SCNS training process is mathematically equivalent to the iterative decoding approach in LDPC codes over the binary erasure channel (BEC) [14, 15, 16] and Irregular Repetition Slotted ALOHA (IRSA) [17, 18, 19, 20, 21, 22]. Following the density evolution analysis (in the asymptotic regime of a large number of skills) [23, 24, 25], let $q^{(i)}$ be the probability that the skill end of a randomly selected edge is not learned after the $i^{th}$ iteration of training. Initially, we have $q^{(0)}=1$ (as all the skills are novel before training). After the $i^{th}$ iteration, the number of novel skills in a randomly selected text follows a Poisson distribution with mean $q^{(i)}c$ (a more detailed argument for this is provided in Section VII-A). For the 1-skill learner, a novel skill can be learned if it is the only novel skill in a text, which happens with probability $e^{-q^{(i)}c}$. Thus, the probability that the text end of a randomly selected edge is not learned after the $i^{th}$ iteration is
$p^{(i)}=1-e^{-q^{(i)}c}.$ (1)
@@
Therefore, by SCNS, a skill end $s$ of a randomly selected edge can not be learned after the $(i+1)^{th}$ iteration if the text nodes connected to $s$ (excluding the text end of the randomly selected edge) were not able to learn the skill $s$ after the $i^{th}$ iteration. Since the excess degree of a Poisson degree distribution is still Poisson with the same mean [9], the number of text nodes connected to $s$ (excluding the text end of the randomly selected edge) is a Poisson random variable with mean $cR$ (from Proposition 3(i)). Thus,
$q^{(i+1)}=\sum_{k=0}^{\infty}e^{-cR}\frac{(cR)^{k}}{k!}{p^{(i)}}^{k}=e^{-cR(1-p^{(i)})}.$ (2)
@@
The equation in (3) characterizes the scaling law of the value $p$ with respect to $R$, the ratio of the number of training texts to the number of skills.
@@
Now a skill $s$ remains novel after training if it cannot be learned from any text connected to $s$. As the degree of a skill node is Poisson distributed with mean $cR$, this occurs with probability $e^{-cR(1-p)}$. Thus, the probability that a randomly selected skill is learned is $p_{s}=1-e^{-cR(1-p)}$. For testing, we randomly select a text $t$ from $T$ (not included in the $D$ training texts). The number of skills in $t$ is also Poisson distributed with mean $c$. Then the probability that the text is understood by the learner is
$\sum_{k=0}^{\infty}e^{-c}\frac{c^{k}}{k!}(p_{s})^{k}=e^{-c(1-p_{s})}=e^{-ce^{-cR(1-p)}}.$ (4)
@@
By the equality in (3), the testing error for the 1-skill learner (i.e., $1$ minus the probability in (4)) can be expressed as $p$, where $0\leq p\leq 1$ is the largest solution of (3). Since $p\geq 0$, we know from (4) that the testing error is at least $1-e^{-ce^{-cR}}$.
Figure 3: The plot illustrates the computed values of $p$ as a function of $R$, which ranges from 0.1 to 5 with a step size of 0.1, for each $c$ value (1, 2, 3, 4, 5). The computation is based on the recursive equation $p^{(i+1)}=1-e^{-(ce^{-cR(1-p^{(i)})})}$ with the initial condition $p^{(0)}=1$.
@@
In Figure 3, we illustrate the effect of the parameters $R$ and $c$ on the values of $p$. We compute $p$ by using the recursive equation $p^{(i+1)}=1-e^{-(ce^{-cR(1-p^{(i)})})}$ with the initial condition $p^{(0)}=1$. The number of iterations is set to 10,000. Since $1-p$ also represents the probability that a (test) text is understood by the learner, there exists a percolation threshold, indicated by a sudden change in the values of $p$. The phase transition behavior is most evident when the average number of skills contained in each text (i.e., $c$) is greater than $3$. For $c=3$ (respectively, 4 and 5), this threshold occurs when $R$ exceeds approximately 1 (respectively, 1.5 and 2.5). The underlying rationale is that for $c\geq 3$, a randomly chosen text is unlikely to encompass precisely one novel skill. Consequently, a substantial volume of training texts is essential to ensure a sufficient quantity of texts initially containing only a single novel skill. This foundation of learned skills subsequently initiates a cascade of successive cancellations of novel skills.

### v3

@@ Abstract
Recent advances in Large Language Models (LLMs) have demonstrated the emergence of capabilities (learned skills) when the number of system parameters and the size of training data surpass certain thresholds. The exact mechanisms behind such phenomena are not fully understood and remain a topic of active research. Inspired by the skill-text bipartite graph model proposed by Arora and Goyal for modeling semantic languages, we develop a mathematical theory to explain the emergence of learned skills, taking the learning (or training) process into account. Our approach models the learning process for skills in the skill-text bipartite graph as an iterative decoding process in Low-Density Parity Check (LDPC) codes and Irregular Repetition Slotted ALOHA (IRSA). Using density evolution analysis, we demonstrate the emergence of learned skills when the ratio of the number of training texts to the number of skills exceeds a certain threshold. Our analysis also yields a scaling law for testing errors relative to this ratio. Upon completion of the training, the association of learned skills can also be acquired to form a skill association graph. We use site percolation analysis to derive the conditions for the existence of a giant component in the skill association graph. Our analysis can also be extended to the setting with a hierarchy of skills, where a fine-tuned model is built upon a foundation model. It is also applicable to the setting with multiple classes of skills and texts. As an important application, we propose a method for semantic compression and discuss its connections to semantic communication.
Index Terms:
Large language models, emergence of capabilities, Low-Density Parity Check codes, Irregular Repetition Slotted ALOHA, density evolution, semantic communication
@@ I Introduction
In the recent era of natural language processing (NLP), the evolution of large language models (LLMs), such as GPT-4 [1] and Gemini [2], has greatly impacted people’s daily lives. There is a growing consensus that enhancing language model performance and sample efficiency across a broad spectrum of downstream NLP tasks is closely linked to scaling up these models. This scaling involves increasing both the size of the training data and the number of model parameters, as discussed in [3, 4]. The relationship between the scale of an LLM and its performance can often be quantitatively forecasted using scaling laws, as explored in [5, 6].
@@
    Learning as an Iterative Decoding Process: We model the learning/training process for the skills in the skill-text bipartite graph as an iterative decoding process, similar to that adopted in Low-Density Parity Check (LDPC) codes [14, 15, 16] and Irregular Repetition Slotted ALOHA (IRSA) [17, 18, 19, 20, 21, 22]. When a text is presented to a learner, a skill in that text might be learned with a certain probability (specific definitions for learners are given in Section III). By repeatedly presenting a large number of training texts to a learner, a fraction of skills can be learned.
@@
    The Emergence of Learned Skills: Following the density evolution analysis for the iterative decoding process (in the asymptotic regime of a large number of skills) [23, 24, 25], we show the emergence of learned skills when the ratio $R$ of the number of training texts to the number of skills exceeds a certain threshold. Once this threshold is exceeded, the testing error, defined as the probability that a randomly selected text is not understood by the learner, drops sharply. Our analysis also provides the scaling law of the testing error with respect to the ratio $R$.
@@
    The Association of Skills: Upon completing training, a learner acquires not just isolated skills but also the interrelation of these skills, which can be represented by a skill association graph where edges connect learned skills appearing in the same text. Understanding this graph is crucial for making inferences, such as predicting the next skill from a given set. This process is akin to prompting a language model to forecast and generate a text based on a skill set. This leads to a critical inquiry: Are the learned skills sufficiently interconnected to aid in prediction? To address this question, we use site (or node) percolation analysis [9] to derive conditions for the existence of a giant component in the skill association graph.
@@
The paper is organized as follows. In Section II, we define a semantic language using a skill-text bipartite graph. In Section III, we introduce abstract learners, including 1-skill learners in Section III-A and Poisson learners in Section III-B. Density evolution analysis is employed to derive the scaling law for these abstract learners. In Section IV, we address the problem of learning the association between two learned skills and derive conditions for the existence of a giant component in the skill association graph. In Section V, we further address the problem setting with a hierarchy of skills by using a foundation model and a fine-tuning model. We present the extension to the setting with multiple classes of skills and texts in Section VI. In Section VII, we demonstrate how the trained learners can be utilized for semantic compression and semantic communication. The paper concludes in Section VIII, where we discuss possible extensions of our work.
@@ II-A Definition of a Semantic Language
[…]
@@
In this paper, we address the question of learning a semantic language. A common approach is to sample a subset of texts from the set of texts ${\cal T}$ and present this subset to a learner to learn the mapping $\phi$.
Definition 2
A skill $s$ in a semantic language is learned if the learner can determine whether the skill $s$ is present in any given text $t$ in $𝒯\color[rgb]{0,0,0}{\cal T}$. A text $t$ is understood by a learner if all the skills contained within $t$ are learned by the learner.
II-B Learning a Semantic Language by Sampling
Suppose that a subset of texts ${\cal D}$ is sampled from the set of texts $𝒯\color[rgb]{0,0,0}{\cal T}$ and presented to the learner as training texts. Let $𝒮R=|{\cal D}|/|\color[rgb]{0,0,0}{\cal S}|$ be the ratio of the number of training texts to the number of skills.
We make the following two assumptions about the sampled texts:
(Poisson degree distribution) Let $N(t)=|\phi(t)|$ be the number of skills in the text $t\in{\cal D}$. Then $\{N(t),t\in{\cal D}\}$ are independent Poisson random variables with mean $c$, i.e., ${\bf\sf P}(N(t)=k)=e^{-c}\frac{c^{k}}{k!}$ for $k=0,1,2,\ldots$.
(Uniform connection) Each edge of a text node is connected to a skill node independently and uniformly. Thus, the probability that an edge of the text node $t$ is connected to a particular skill node $s$ is $1/|{\cal S}|$.
@@ II-B Learning a Semantic Language by Sampling
From (A1) and (A2), the skill-text bipartite graph of a sampled semantic language is a random bipartite graph with $|{\cal S}|$ skill nodes on one side and $|{\cal D}|$ text nodes on the other side. Denote by $M(s,t)$ the number of edges between a skill node $s$ and a text node $t$. Since randomly splitting a Poisson process yields independent Poisson processes, we have from (A1) and (A2) that $\{M(s,t),s\in{\cal S},t\in{\cal D}\}$ are independent Poisson random variables with mean $c/|{\cal S}|$. Denote by $M(s)$ the number of edges of a skill node $s$. As a superposition of independent Poisson processes (from the $|{\cal D}|$ text nodes) is a Poisson process, we know that $\{M(s),s\in{\cal S}\}$ are independent Poisson random variables with mean $c|{\cal D}|/|{\cal S}|$ (i.e., $cR$). On the other hand, the probability that a point in a superposition of independent Poisson processes comes from one of these processes is proportional to their rates. This leads to the following proposition.
@@
Consider a sampled skill-text bipartite graph from a semantic language ${\cal L}=({\cal A},{{\cal T},{\cal S},}\phi)$. Then:
(Poisson degree distribution) $\{M(s),s\in{\cal S}\}$ are independent Poisson random variables with mean $c|{\cal D}|/|{\cal S}|=cR$, where $M(s)$ is the degree of the skill node $s$.
(Uniform connection) Each edge of a skill node is connected to a text node independently and uniformly. Thus, the probability that an edge of a skill node is connected to a particular text node is $1/|{\cal D}|$.
Therefore, such a bipartite graph is a random bipartite graph, as per the configuration model in [9], where the degree distribution of the text nodes (respectively, skill nodes) is Poisson with mean $c$ (respectively, $cR$).
@@ III Abstract Learners
In this section, we develop a mathematical theory to understand the learning process of a semantic language by an abstract learner. To a learner, all skills are novel before training, but a skill $s$ in text $t$ may be learned after the text is presented to the learner. By repeatedly presenting a set of training texts ${\cal D}$, we aim to determine the fraction of skills that can be learned by the learner. We consider two types of learners based on their learning capabilities, namely, a 1-skill learner and a more general Poisson learner.
@@
(1-Skill Learner) An abstract learner is called a 1-skill learner if it can learn a novel skill $s$ by being presented with a text $t$ where the skill $s$ is the only novel skill in the text. Once a skill $s$ is learned, the learner is able to identify the skill $s$ appearing in any other text.
@@
Now we describe how the 1-skill learner is trained. For each iteration, we present the $|{\cal D}|$ sampled texts to the 1-skill learner in parallel. In the first iteration, a text containing only one skill is used to learn the skill in that text. Texts containing more than one skill do not contribute to learning in the first iteration. Once a skill $s$ is learned by the 1-skill learner, the number of novel skills in other texts containing $s$ is reduced by 1. In other words, the edges connected to the skill node $s$ can be removed from the skill-text bipartite graph. Therefore, the number of novel skills associated with each text in the next iteration will not be greater than that in the current iteration. In the second iteration, texts with only one novel skill remaining are used to learn the skills in those texts. As in the first iteration, skills learned in the second iteration can be used to remove the corresponding edges in the skill-text bipartite graph. This iteration process is repeated until no more novel skills can be learned. We refer to this training process as the Successive Cancellation of Novel Skills (SCNS).
Figure 2: An example of the Successive Cancellation of Novel Skills (SCNS).
@@
In Figure 2, we present a simple example of SCNS. Figure 2(a) depicts the original skill-text bipartite graph. Figure 2(b) focuses on the skill node $a$, which is connected to a text node with degree $1$; hence, we use a red edge to indicate that the skill node $a$ is learned by the text node. Consequently, the skill node $a$ can be learned, allowing the edges connected to it to be removed from the skill-text bipartite graph. Figure 2(c) shows the graph after the edges are removed, revealing that the skill node $b$ is connected to a text node with degree $1$, indicating it can also be learned. Figure 2(d) represents the last iteration, where no additional skills can be learned, concluding the process with the skills $a$ and $b$ being learned, while skills $c$ and $d$ are not.
@@
This SCNS training process is mathematically equivalent to the iterative decoding approach in LDPC codes over the binary erasure channel (BEC) [14, 15, 16] and Irregular Repetition Slotted ALOHA (IRSA) [17, 18, 19, 20, 21, 22]. Following the density evolution analysis (in the asymptotic regime of a large number of skills) [23, 24, 25], we can derive the probability that a randomly selected text is not understood by the learner. This can be viewed as the training error of the learner as a result of the above training process. In particular, let $q^{(i)}$ be the probability that the skill end of a randomly selected edge is not learned after the $i^{th}$ SCNS iteration of training. Initially, we have $q^{(0)}=1$ (as all the skills are novel before training). After the $i^{th}$ SCNS iteration, the number of novel skills in a randomly selected text follows a Poisson distribution with mean $q^{(i)}c$ (a more detailed argument for this is provided in Section VI-A). For the 1-skill learner, a novel skill can be learned if it is the only novel skill in a text, which happens with probability $e^{-q^{(i)}c}$. Thus, the probability that the text end of a randomly selected edge does not contribute to the learning of a novel skill in the $(i+1)^{th}$ SCNS iteration is
$p^{(i+1)}=1-e^{-q^{(i)}c}.$ (1)
@@
To compute $q^{(i+1)}$, let us consider a randomly selected edge. Suppose that the skill (resp. text) end of that edge is $s$ (resp. $t$). By SCNS, the skill $s$ is not learned after the $(i+1)^{th}$ SCNS iteration if the text nodes connected to $s$ (excluding the text node $t$ from the randomly selected edge) are not able to contribute to the learning of skill $s$ in the $(i+1)^{th}$ SCNS iteration. Since the excess degree of a Poisson degree distribution is still Poisson with the same mean [9], the number of text nodes connected to $s$ (excluding the text node $t$) is a Poisson random variable with mean $cR$ (from Proposition 3(i)). Thus,
$q^{(i+1)}=\sum_{k=0}^{\infty}e^{-cR}\frac{(cR)^{k}}{k!}{p^{(i+1)}}^{k}=e^{-cR(1-p^{(i+1)})}.$ (2)
@@
Here, $p$ represents the probability that the text end of a randomly selected edge still contains more than one novel skill at the end of the training process (i.e., after the learner has been repeatedly presented with the set of training texts ${\cal D}$). Equation (3) characterizes the scaling law of the value $p$ with respect to $R$, the ratio of the number of training texts to the number of skills.
@@
Now a skill $s$ remains novel after training if it cannot be learned from any text connected to $s$. (see Figure 3 for an illustration). As the degree of a skill node is Poisson distributed with mean $cR$, this occurs with probability $e^{-cR(1-p)}$. Thus, the probability that a randomly selected skill is learned is
$\zeta=1-e^{-cR(1-p)}.$ (4)
Figure 3: We illustrate the density evolution (tree) analysis by considering a skill node $s$. The skill node $s$ remains novel if it cannot be learned from any text connected to $s$. This leads to (4) (as the degree of node $s$ is Poisson with mean $cR$). For the 1-skill learner, the text end $t$ of an edge connected to $s$ can learn $s$ if it is the only edge connected to $t$, i.e., all the other edges connected to $t$ (excluding the edge between $s$ and $t$) are removed. This leads to (1) (as the excess degree of node $t$, excluding the edge between $s$ and $t$, is Poisson with mean $c$and each excess edge is removed with probability $q$). Now let us go down the tree one level further. The edge between $t$ and $s^{\prime}$ cannot be removed if $s^{\prime}$ is not learned (remains novel). The skill node $s^{\prime}$ remains novel if it cannot be learned from any text connected to $s^{\prime}$ (excluding the text $t$). This leads to (2) (as the excess degree of node $s^{\prime}$, excluding the edge between $s^{\prime}$ and $t$, is Poisson with mean $cR$).
For testing, we randomly select a text $t$ from ${\cal T}$ (not included in the set of training texts ${\cal D}$). Assume that this randomly selected text has the same distribution as those in the training data, i.e., it also satisfies (A1) and (A2). As such, the number of skills in $t$ is also Poisson distributed with mean $c$. Recall that a text can be understood if all the connected skill nodes are learned (see Figure 4 for an illustration). Define the testing error (denoted by $\epsilon$) as the probability that the randomly selected text is not understood by the learner. For the 1-skill learner, we have
$\epsilon$ $=$ $1-\sum_{k=0}^{\infty}e^{-c}\frac{c^{k}}{k!}\zeta^{k}$ (5)
$=$ $1-e^{-c(1-\zeta)}=1-e^{-ce^{-cR(1-p)}}.$
@@
In this case, the probability of the testing error is also equal to the solution for $p$ in (3), for $0\leq p\leq 1$. Since $p\geq 0$, we know from (5) that the testing error is at least $1-e^{-ce^{-cR}}$.
Figure 4: A randomly selected text can be understood if all the connected skill nodes are learned (highlighted in red circles). This leads to (5).
Figure 5: The plot illustrates the probability of the testing error $\epsilon$ as a function of $R$ for the 1-skill learner.
@@
In Figure 5, we illustrate the effect of the parameters $R$ and $c$ on the probability of the testing error $\epsilon$. We compute $p$ by using the recursive equation $p^{(i+1)}=1-e^{-ce^{-cR(1-p^{(i)})}}$ with the initial condition $p^{(0)}=1$. The number of iterations is set to 10,000. We can see that there exists a percolation threshold, indicated by a sudden change in the values of $\epsilon$. The phase transition behavior is most evident when the average number of skills contained in each text (i.e., $c$) is greater than $3$. For $c=3$ (respectively, 4 and 5), this threshold occurs when $R$ exceeds approximately 1 (respectively, 1.5 and 2.5). The underlying rationale is that for $c\geq 3$, a randomly chosen text is unlikely to encompass precisely one novel skill. Consequently, a substantial volume of training texts is essential to ensure a sufficient quantity of texts initially containing only a single novel skill. This foundation of learned skills subsequently initiates a cascade of successive cancellations of novel skills.